OpenStudy (anonymous):

Use the fact that x2 + 10x + 25 = (x + 5)2 to find the expression. √x^2+10x+25

OpenStudy (anonymous):

i really have no clue what they are asking me to do or find out. can anybody help?

OpenStudy (asnaseer):

\[(x+5)^2=x^2+10x+25\]so it follows that:\[\sqrt{x^2+10x+25}=\sqrt{(x+5)^2}=x+5\]or, strictly speaking:\[\pm(x+5)\]